OpenStudy (anonymous):

Find the missing measure for a right circular cone given the following information. Find V if r = 2 and L.A. = 12pi

OpenStudy (anonymous):

so \[12 \pi = \pi r \sqrt{h^2 + r^2}\]and\[V=\frac{ 1 }{ 3 }\pi r^2 h\]right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

and r = 2?

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

so you can solve the first equation for h, correct? then you can use that value of h with r = 2 in the second equation to solve for the volume...

OpenStudy (anonymous):

but you need the slant height as well to solve this

OpenStudy (anonymous):

the slant height is \(\sqrt{h^2+r^2}\) where h is the vertical height and r is the radius.

OpenStudy (anonymous):

oh okay i get it now thanks :)

OpenStudy (anonymous):

you're welcome!